Use two equations in two variables to solve the application.
The sum of two integers is 45, and their difference is 13. Find the integers (smaller and larger).

Answer by addingup(3677)

(1) x + y = 45
(2) x - y = 13 therefore x = y + 13 Substitute for x on (1)
y + 13 + y = 45
2y = 32
y = 16
and x - y = 13; x - 16 = 13; x = 29
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
x + y = 45
29 + 16 = 45 Correct